OpenStudy (anonymous):

Calculate the kinetic energy, in joules, of the electrons ejected from the surface of cesium by radiation of wavelength (4.20x10^2) nm. Take the threshold frequency of copper to be vo = 4.38x1014 Hz. I have no idea how to even go about this...

OpenStudy (anonymous):

vo = threshold frequency.

OpenStudy (anonymous):

ok you first need to transfer wavelenght to frequency to se weather your radiation has enough energy to eject electrons (if value is greater than treshold) and then if it is use E = h* ni where h - planck constant ni - frequency

OpenStudy (anonymous):

Okay, when I calculated frequency I got \[7.138095\times10^{14}\] s-1 I inserted that into e = h * ni and got \[4.7297019\times10^{-19}\] Joules. Does that seem right?

OpenStudy (anonymous):

ok if you put numbers corectly in calculator you got that frequency is greather than threshold and that is good it means you can continue with calculation of energy, and then if you calculated it corectly it should be right

OpenStudy (anonymous):

I got it! Thanks!

OpenStudy (anonymous):

not a problem, glad to help